How many 6-digit numbers are there in 4,6,2,3,7 and 9 if repetition is not allowed?
Answer by Theo(13342)

there would be 6! = 6*5*4*3*2*1 = 720
if only 2 numbers (4 and 6), then the number of 2 digit numbers would be equal to 2! = 2 * 1 = 2 and the numbers would be 46, 64.
if only 3 numbers (4 and 6 2), then the number of 3 digit numbers would be equal to 3! = 3 * 2 * 1 = 6 and the numbers would be 462, 426, 642, 624, 246, 264.
720 numbers with 6 digits is much harder to show, but you can see that the formula applies to any number of digits.